# Long Context Scaling: Divide and Conquer via Multi-Agent Question-driven Collaboration

### v1

@@ Abstract
Processing long contexts has become a critical capability for modern large language models (LLMs). Existing works leverage agent-based divide-and-conquer methods for processing long contexts. But these methods face crucial limitations, including prohibitive accumulated latency and amplified information loss from excessive agent invocations, and the disruption of inherent textual dependencies by immoderate partitioning. In this paper, we propose a novel multi-agent framework XpandA (Expand-Agent) coupled with question-driven workflow and dynamic partitioning for robust long-context processing. XpandA overcomes these limitations through: 1) dynamic partitioning of long texts, which adaptively modulates the filling rate of context windows for input sequences of vastly varying lengths; 2) question-guided protocol to update flat information ensembles within centralized shared memory, constructing consistent inter-agent knowledge across partitions; and 3) selectively replaying specific partitions based on the state-tracking of question-information couples to promote the resolution of inverted-order structures across partitions (e.g., flashbacks). We perform a comprehensive evaluation of XpandA on multiple long-context benchmarks with length varying from 1k to 1M, demonstrating XpandA’s feasibility for processing ultra-long sequences and its significant effectiveness in enhancing the long-context capabilities of various LLMs by achieving 20% improvements and 1.5x inference speedup over baselines of full-context, RAG and previous agent-based methods.
@@ 1 Introduction
Like human cognition—which depends on long-term memory, accumulated knowledge, and life experience—Artificial General Intelligence (AGI) development requires the ability to associate distant information and retain context over extended content [38; 58]. Long-context tasks, such as question answering [76; 92; 68], summarization [105] and coding [7], consistently present significant challenges due to issues like hallucination and overlooking critical information [50].
Recent efforts to tackle these challenges have focused on three key directions: training data strategy, architecture and workflow design. The training data strategy conducts long-context data filtering [86; 71], mixture [23; 33], and synthesis [24; 27] on pre-training and post-training phase. Architecture methods employ position embeddings [73; 64; 25; 74] and attention mechanisms such as transformer-based [5; 88; 108; 95; 22], linear-complexity [65] and hybrid architectures [59; 47; 53]. Despite these efforts, the lack of extensibility in context windows leads to input truncation when processing longer sequences, while training long-context language models entails prohibitively high computational costs[5]. Workflow design, including external memory modules [80; 56], retrieval-augmented generation (RAG) [8; 34], and agent-based methods (LongAgent [101], Chain of Agents [100]), etc., can mitigate the challenges mentioned above but introduce issues such as information loss for inappropriate extraction [100], Inter-Agent Misalignment and Specification Violation inherent in the multi-agent system [9]. Furthermore, the most concerning aspect of agent-based methods is the frequent model invocation. This not only potentially amplifies intermediate errors but also leads to an accumulation of model call latency.
@@ 2.1 Long Context Language Modeling
Long context processing in LLMs has recently become a central focus [49], driven by tasks such as long-document QA [35; 15; 76; 55; 21], summarization [29; 105], retrieval-augmented generation (RAG) [39], in-context learning [46; 90], and large-codebase understanding [32; 7]. These require models to locate, aggregate, and reason over dispersed information, while overcoming "lost in the middle"[50] issues. Existing approaches fall into three main categories: Training data strategies, architectural and workflow design [48]. Data strategies enhance a model’s capacity to process lengthy texts through the filter [86; 71], mixture [23; 33], and synthesis [24; 27] of long-context data in both pre-training and post-training phases. Architecture design aims to develop model structures capable of effectively handling extended texts by position embeddings [73; 64; 25; 74] and attention mechanisms such as transformer-based [5; 88; 108; 95; 22], linear-complexity [65] and hybrid architectures [59; 47; 53]. Workflow-based methods minimize model changes, using techniques like external memory [80; 56], RAG [8; 34], or agent-based reasoning [11; 45; 100], to further improve the long context capability.
@@ 2.2 LLM Reasoning with Search and Planning
Human problem-solving relies on decomposing complex tasks into manageable steps [19], a paradigm successfully adapted to LLMs. Early work by [85] and [36] introduced Chain-of-Thought (CoT) prompting, using intermediate steps to guide reasoning, while later studies like [20] and [106] proposed iterative decomposition into subtasks. To enhance reasoning structures, [107] and [3] developed frameworks for self-discovering and dynamically adapting reasoning modules. Further, methods like [17] and [99] address error correction in intermediate steps, particularly for smaller models. Beyond linear reasoning, exploration-based approaches leverage search over multiple paths, inspired by human problem-solving [72]. While early methods like self-consistency [81] generated limited diversity, advanced frameworks such as Tree-of-Thoughts [93] and Graph-of-Thoughts [6] enable finer-grained branching. Aggregation strategies further improve robustness, ranging from ensemble-based voting [77; 44] to verifier-guided selection [82].
@@ 2.3 Multi-Agent LLMs
Research efforts in Multi-Agent Systems have encompassed diverse aspects of collaboration, covering mechanisms like cooperation [14; 70], competition [12; 102], and coopetition [1; 16], as well as strategies such as rule-based [97; 109], role-based [14; 28], and model-based approaches [42; 89]. Furthermore, other key considerations include communication structures (centralized [31; 61], decentralized [14; 94], and hierarchical [41; 52]), which offer scenario-dependent advantages and disadvantages, and coordination architectures (static or dynamic [13; 30; 83]), providing trade-offs between stability and flexibility. MASs demonstrate broad potential for applications across numerous domains, including 5G/6G networks [84; 104], Natural Language Generation (NLG) and Question Answering (QA) [31; 75; 87] and socio-cultural simulations [40; 60], among others. Nevertheless, the field still confronts numerous challenges [10]. These include achieving unified governance, understanding the inherent limitations of individual agents [97; 69], scalability, efficient resource management [79], developing robust evaluation benchmarks [51; 63] and addressing ethical risks and safety concerns [2; 18; 57].
Input: Context $C$, Query $Q$.
Parameters: LLM$(\theta)$, instruct prompts $(I_{E},I_{D})$, chunk size $K$, chunk overlap $P$, max replay times MRT
Output: Answer $A$.
Split $C$ into $x$ chunks $\{c_{1},c_{2},\cdots,c_{x}\}$;
Create empty dictionary $T$, $P$;
Initialize $o\leftarrow 1$, $rev\leftarrow 1$;
@@
       for $i$ in $o,o+rev,\cdots$ do
             $T_{i},P_{i}\leftarrow\text{Explorer}(I_{E},Q,T,P,c_{i})$;
            $T\leftarrow\text{Merge}(T,T_{i})$;
             $P\leftarrow\text{Merge}(P,P_{i})$;
            for every $k\in P_{i}.\text{keys}$ and $k\in T.\text{keys}$ do
                   $\text{Delete}(k,T[k])$ from $T$;
             end for
       end for
      $\text{Action},\hat{A}\leftarrow\text{Decider}(I_{D},Q,P)$;
       if Replay in Action then
             $o\leftarrow\begin{cases}\max(\min(T.values-1,1)&\text{if }rev>0\\ \min(\max(T.values+1,l)&\text{otherwise}\end{cases}$
            $rev\leftarrow-rev$;
@@
       end if
      else Break;
end while
$A\leftarrow\text{Parse}(\hat{A})$;
return $A$;
Algorithm 1 XpandA Workflow
3 Method
@@ 3.1 Stage 1: Dynamic Partition of Long Input Sequence
Chunk $\text{Chunk size}=\begin{cases}\left\lceil\frac{w}{n}\right\rceil&\text{if }\frac{w}{n}\leq M\\ M&\text{if }\frac{w}{n}>M\end{cases}$ (1)
[…]
Number of $\text{Number of chunks}=\left\lceil\frac{w}{M-\delta}\right\rceil$ (2)
@@ 3.2 Stage 2: Question-guided Exploration
In Stage 1: XpandA operates on a sequence of $l$ Explorers. Each Explorer accepts the concatenation of a document chunk $c_{i}$, user query $q$, instruction for a specific reasoning task $I_{D}$ and gathered information, denoted as "(question,[answer]) pairs" $P$. Explorer is instructed to output enlarged $P_{i}$ and generate newly-generated problems included in "Unsolved Problem Tracer" $T_{i}$ expressed as:
@@
$T$,$P$ will be updated with $T_{i}$,$P_{i}$ when the processed chunk is switched.
@@
$Expand$ function has two types of behaviors: 1) Breaking down the generalized problem into sub-problems, and 2) expanding new problem entities based on a problem and the corresponding information. While the loop iterates, the problem-solving focus of large language model is gradually expanded across the chunks.
@@ 3.3 Stage 3: Iterative Selective Replay
This dependency forms an AOV network where chunks are vertices and dependencies are edges. When MRT (max replay times) $>x-1$, XpandA guarantees to complete dependency resolution A if LLM can finish information extraction in a chunk unit.
4 Experiment
@@ 4.1 Experiment Setup
[…]
  • Question Answering. HotpotQA [92] is a Wikipedia-based multihop QA dataset. It requires reasoning across multiple passages to find the answer. NarrativeQA [68] is a dataset focusing on answering questions based on stories or scripts, including understanding of important elements such as characters, plots, themes, etc. loogle-MR-mixup and loogle-CR-mixup originate from LooGLE [43] Long-dependency QA task, specifically the Multiple information Retrieval and Comprehension and Reasoning subtasks and are mixed up with mix up distracting documents and supporting documents of stepwise growing length.
  • Retrieval. OpenAI-MRCR is a long context dataset for benchmarking an LLM’s ability to distinguish between multiple needles hidden in context. The evaluation consists of retrieval of the response corresponding to a specific instance from multi-turn synthetic dialogue.
[…]
LLM Backbones. The backbone models we deploy through include the well-established and popular open-source model Llama3.1-8B-Instuct [26], Qwen2.5-7B-Instruct [66] as well as the latest long-context language model Qwen2.5-7B-Instruct-1M [91] model. We choose efficient engine vllm [37] and 4*A100_80G GPU as infrastructures for the experiments to handle immense GPU memory required for ultra-long input sequences up to 1M tokens.
Baselines. Full-context directly places the query and the long text into the context window where truncation may occur. RAG is implemented through RAG-Fusion [67] by utilizing reciprocal rank fusion of relevance and word frequency ranking to reach strong and robust retrieval performance. We introduce previous Agent-based works CoA [100] and LongAgent [101] to enhance the comprehensiveness of our baseline to highlight the improvement of our works.
Metrics. We evaluate the QA tasks and Misc by F1 Score [4] and Sequence Match Ratio [^5] https://docs.python.org/3/library/difflib.html for Retrieval, code tasks by exact match. Besides, we introduce fine-grained Progress Score $f$ from AgentBoard [54] for quantification of the multi-turn LLM agents’ performance which is denoted by:
@@
4.2 Overall Results of XpandA
Table 2: Overall Performance evaluation of XpandA on benchmarks for Question-Answering, Code, Miscellaneous, and Retrieval, showing XpandA achieved optimal or leading Performance over full-context, RAG and previous agent-based works across multiple domains (Wilcoxon Signed-Rank Test: p<0.05, details in Appendix LABEL:apd:st).
Question Answering Code Tasks Misc. Retrieval
@@ 4.2 Overall Results of XpandA
Code Tasks, Retrieval & Misc. Table 2 displays the test results for these three classification tasks. Similarly, XpandA achieved optimal scores on the evaluation metrics for the majority of these tasks. However, it is noteworthy that RAG + Llama3.1-8B-Instruct delivered the best performance on the GraphWalks dataset, with XpandA ranking second. This outcome may be attributed to the fact that the text in GraphWalks is presented in a structured "node->node" format (where each ’node’ is a hexadecimal hash string). Such a structure potentially allows RAG to retrieve the intended textual information with greater precision compared to attention-based language models.

### v3

@@ Abstract
Processing long contexts has become a critical capability for modern large language models (LLMs). Existing works leverage agent-based divide-and-conquer methods for processing long contexts. But these methods face crucial limitations, including prohibitive accumulated latency and amplified information loss from excessive agent invocations, and the disruption of inherent textual dependencies by immoderate partitioning. XpandA tackles long-text challenges using: 1) dynamic partitioning for flexible context handling; 2) question-driven updates for consistent knowledge across partitions; and 3) selective replay of partitions to resolve complex information order (e.g., flashbacks). We perform a comprehensive evaluation of XpandA on multiple long-context benchmarks with length varying from 1k to 1M, demonstrating XpandA’s feasibility for processing ultra-long sequences and its significant effectiveness in enhancing the long-context capabilities of various LLMs by achieving 20% improvements and 1.5x inference speedup over baselines of full-context, RAG and previous agent-based methods.
Machine Learning, ICML
@@ 1 Introduction
Like human cognition—which depends on long-term memory, accumulated knowledge, and life experience—Artificial General Intelligence (AGI) development requires the ability to associate distant information and retain context over extended content (Lake et al., 2016; Morris et al., 2024). Long-context tasks, such as question answering (Wang et al., 2025; Yang et al., 2018; s Koˇciský et al., 2018), summarization (Zhong et al., 2021) and coding (Bogomolov et al., 2024), consistently present significant challenges due to issues like hallucination and overlooking critical information (Liu et al., 2023).
Recent efforts to tackle these challenges have focused on three key directions: training data strategy, architecture and workflow design. The training data strategy conducts long-context data filtering (Wu et al., 2025; Si et al., 2025), mixture (Fu et al., 2024b; Jin et al., 2023), and synthesis (Gao et al., 2025; He et al., 2024) on pre-training and post-training phase. Architecture methods employ position embeddings (Su et al., 2021; Press et al., 2021; Golovneva et al., 2024; Sun et al., 2022) and attention mechanisms such as transformer-based (Beltagy et al., 2020; Xiao et al., 2024; Zhou et al., 2025; Yuan et al., 2025; Fu et al., 2024a), linear-complexity (Qin et al., 2024) and hybrid architectures (Munkhdalai et al., 2024; Lieber et al., 2024; Lu et al., 2025). Despite these efforts, the lack of extensibility in context windows leads to input truncation when processing longer sequences, while training long-context language models entails prohibitively high computational costs(Beltagy et al., 2020). Workflow design, including external memory modules (Wang et al., 2023b; Mehta et al., 2023), retrieval-augmented generation (RAG) (Borgeaud et al., 2022; Joshi et al., 2024), and agent-based methods (LongAgent (Zhao et al., 2024a), Chain of Agents (Zhang et al., 2024d)), etc., can mitigate the challenges mentioned above but introduce issues such as information loss for inappropriate extraction (Zhang et al., 2024d), Inter-Agent Misalignment and Specification Violation inherent in the multi-agent system (Cemri et al., 2025b). Furthermore, the most concerning aspect of agent-based methods is the frequent model invocation. This not only potentially amplifies intermediate errors but also leads to an accumulation of model call latency.
@@ 2.1 Long Context Language Modeling
Long context processing in LLMs has recently become a central focus (Liu et al., 2025b), driven by tasks such as long-document QA (Kociský et al., 2017; Dasigi et al., 2021; Wang et al., 2025; Masry & Hajian, 2024; Fan et al., 2024), summarization (Huang et al., 2021; Zhong et al., 2021), retrieval-augmented generation (RAG) (Lee et al., 2024), in-context learning (Li et al., 2024e; Xu et al., 2024), and large-codebase understanding (Jimenez et al., 2024; Bogomolov et al., 2024). These require models to locate, aggregate, and reason over dispersed information, while overcoming ”lost in the middle”(Liu et al., 2023) issues. Existing approaches fall into three main categories: Training data strategies, architectural and workflow design (Liu et al., 2025a). Data strategies enhance a model’s capacity to process lengthy texts through the filter (Wu et al., 2025; Si et al., 2025), mixture (Fu et al., 2024b; Jin et al., 2023), and synthesis (Gao et al., 2025; He et al., 2024) of long-context data in both pre-training and post-training phases. Architecture design aims to develop model structures capable of effectively handling extended texts by position embeddings (Su et al., 2021; Press et al., 2021; Golovneva et al., 2024; Sun et al., 2022) and attention mechanisms such as transformer-based (Beltagy et al., 2020; Xiao et al., 2024; Zhou et al., 2025; Yuan et al., 2025; Fu et al., 2024a), linear-complexity (Qin et al., 2024) and hybrid architectures (Munkhdalai et al., 2024; Lieber et al., 2024; Lu et al., 2025). Workflow-based methods minimize model changes, using techniques like external memory (Wang et al., 2023b; Mehta et al., 2023), RAG (Borgeaud et al., 2022; Joshi et al., 2024), or agent-based reasoning (Chen et al., 2023; Li et al., 2024d; Zhang et al., 2024d), to further improve the long context capability.
@@ 2.2 LLM Reasoning with Search and Planning
Human problem-solving relies on decomposing complex tasks into manageable steps (Dostál, 2015), a paradigm successfully adapted to LLMs. Early work by (Wei et al., 2022) and (Kojima et al., 2022) introduced Chain-of-Thought (CoT) prompting, using intermediate steps to guide reasoning, while later studies like (Dua et al., 2022) and (Zhou et al., 2023) proposed iterative decomposition into subtasks. To enhance reasoning structures, (Zhou et al., 2024) and (Aswani et al., 2024) developed frameworks for self-discovering and dynamically adapting reasoning modules. Further, methods like (Deng et al., 2024) and (Zhang et al., 2024c) address error correction in intermediate steps, particularly for smaller models. Beyond linear reasoning, exploration-based approaches leverage search over multiple paths, inspired by human problem-solving (Stanovich & West, 2000). While early methods like self-consistency (Wang et al., 2023c) generated limited diversity, advanced frameworks such as Tree-of-Thoughts (Yao et al., 2023) and Graph-of-Thoughts (Besta et al., 2024) enable finer-grained branching. Aggregation strategies further improve robustness, ranging from ensemble-based voting (Wang et al., 2024a; Li et al., 2024c) to verifier-guided selection (Wang & Zhou, 2024).
@@ 2.3 Multi-Agent LLMs
Research efforts in Multi-Agent Systems have encompassed diverse aspects of collaboration, covering mechanisms like cooperation (Chen et al., 2024c; Shinn et al., 2023), competition (Chen et al., 2024a; Zhao et al., 2024b), and coopetition (Abdelnabi et al., 2024; Davidson et al., 2024), as well as strategies such as rule-based (Zhang et al., 2024a; Zhuang et al., 2024), role-based (Chen et al., 2024c; Hong et al., 2024), and model-based approaches (Li et al., 2023b; Xu et al., 2023). Furthermore, other key considerations include communication structures (centralized (Jiang et al., 2023; Ning et al., 2024), decentralized (Chen et al., 2024c; Yin et al., 2023), and hierarchical (Li et al., 2023a; Liu et al., 2024b)), which offer scenario-dependent advantages and disadvantages, and coordination architectures (static or dynamic (Chen et al., 2024b; Jeyakumar et al., 2024; Wang et al., 2024c)), providing trade-offs between stability and flexibility. MASs demonstrate broad potential for applications across numerous domains, including 5G/6G networks (Wang et al., 2024d; Zhao et al., 2024d), Natural Language Generation (NLG) and Question Answering (QA) (Jiang et al., 2023; Suzgun & Kalai, 2024; Wu et al., 2024) and socio-cultural simulations (Li et al., 2024a; Nguyen et al., 2024), among others. Nevertheless, the field still confronts numerous challenges (Cemri et al., 2025a). These include achieving unified governance, understanding the inherent limitations of individual agents (Zhang et al., 2024a; Shayegani et al., 2023), scalability, efficient resource management (Wang et al., 2024b), developing robust evaluation benchmarks (Liu et al., 2024a; Peng et al., 2024) and addressing ethical risks and safety concerns (Akbulut et al., 2024; Deshpande et al., 2023; Meinke et al., 2024).
3 Method
Algorithm 1 XpandA Workflow
1:Input: context $C$, query $Q$
@@ 3 Method
3: Split $C$ into $x$ chunks $\{c_{1},c_{2},\ldots,c_{x}\}$
4: Create empty dictionaries $T,P$
5: Initialize $o\leftarrow 1$, $rev\leftarrow 1$, $\textit{replayCount}\leftarrow 0$
6:repeat
7:  for $i$ in sequence $o,o+rev,o+2rev,\ldots$ (within chunk bounds $1..x$) do
8:   $T_{i},P_{i}\leftarrow\text{Explorer}(I_{E},Q,T,P,c_{i})$
9:   $T\leftarrow\text{Merge}(T,T_{i})$
10:   $P\leftarrow\text{Merge}(P,P_{i})$
11:   for every key $k$ such that $k\in\text{keys}(P_{i})$ and $k\in\text{keys}(T)$ do
12:    Delete entry with key $k$ from $T$
13:   end for
14:  end for
15:  $\text{Action},\hat{A}\leftarrow\text{Decider}(I_{D},Q,P)$
16:  if $\text{"Replay"}\in\text{Action}$ then
17:   $o\leftarrow\begin{cases}\max((\min\text{values}(T))-1,1)&\text{if }rev>0\\ \min((\max\text{values}(T))+1,x)&\text{otherwise}\end{cases}$
18:   $rev\leftarrow-rev$
19:   $\textit{replayCount}\leftarrow\textit{replayCount}+1$
20:  else
21:   Break
22:  end if
23:until $\textit{replayCount}>$ MRT
@@
25:return AA
@@ 3.1 Stage 1: Dynamic Partition of Long Input Sequence
$\text{Chunk size}=\begin{cases}\left\lceil\frac{w}{n}\right\rceil&\text{if }\frac{w}{n}\leq M\\ M&\text{if }\frac{w}{n}>M\end{cases}$ (1)
where $\left\lceil\cdot\right\rceil$ denotes the ceiling function. In the first case, chunks maintain approximately equal size with an overlap $\delta$. In the second case, the number of chunks grows dynamically as:
$\text{Number of chunks}=\left\lceil\frac{w}{M-\delta}\right\rceil$ (2)
@@ 3.2 Stage 2: Question-guided Exploration
In Stage 1: XpandA operates on a sequence of $l$ Explorers. Each Explorer accepts the concatenation of a document chunk $c_{i}$, user query $q$, instruction for a specific reasoning task $I_{D}$ and gathered information, denoted as ”(question,[answer]) pairs” $P$. Explorer is instructed to output enlarged $P_{i}$ and generate newly-generated problems included in ”Unsolved Problem Tracer” $T_{i}$ expressed as:
@@
$T$,$P$ will be updated with $T_{i}$$P_{i}$ when the processed chunk is switched.
@@
ExpandExpand function has two types of behaviors: 1) Breaking down the generalized problem into sub-problems, and 2) expanding new problem entities based on a problem and the corresponding information. While the loop iterates, the problem-solving focus of large language model is gradually expanded across the chunks.
@@ 3.3 Stage 3: Iterative Selective Replay
This dependency forms an AOV network where chunks are vertices and dependencies are edges. When MRT (max replay times) >x1>x-1, XpandA guarantees to complete dependency resolution A if LLM can finish information extraction in a chunk unit.
4 Experiments
@@ 4.1 Experiment Setup
Table 2: Overall Performance evaluation of XpandA on benchmarks for Question-Answering, Code, Miscellaneous, and Retrieval, showing XpandA achieved optimal or leading Performance over full-context, RAG and previous agent-based works across multiple domains. The darker blue areas indicate the globally optimal performance, while the lighter blue and red colors represent the local superiority and inferiority respectively.
Question Answering Code Tasks Misc Retrieval
@@
[…]
  • Question Answering. HotpotQA (Yang et al., 2018) is a Wikipedia-based multihop QA dataset. It requires reasoning across multiple passages to find the answer. NarrativeQA (s Koˇciský et al., 2018) is a dataset focusing on answering questions based on stories or scripts, including understanding of important elements such as characters, plots, themes, etc. loogle-MR-mixup and loogle-CR-mixup originate from LooGLE (Li et al., 2024b) Long-dependency QA task, specifically the Multiple information Retrieval and Comprehension and Reasoning subtasks and are mixed up with mix up distracting documents and supporting documents of stepwise growing length.
  • Retrieval. OpenAI-MRCR (Vodrahalli et al., 2024)is a long context dataset for benchmarking an LLM’s ability to distinguish between multiple needles hidden in context. The evaluation consists of retrieval of the response corresponding to a specific instance from multi-turn synthetic dialogue.
[…]
LLM Backbones. The backbone models we deploy through include the well-established and popular open-source model Llama3.1-8B-Instuct (Grattafiori et al., 2024), Qwen2.5-7B-Instruct (Qwen et al., 2025) as well as the latest long-context language model Qwen2.5-7B-Instruct-1M (Yang et al., 2025) model. We choose efficient engine vllm (Kwon et al., 2023) and 4*A100_80G GPU as infrastructures for the experiments to handle immense GPU memory required for ultra-long input sequences up to 1M tokens.
Baselines. Full-context directly places the query and the long text into the context window where truncation may occur. RAG is implemented through RAG-Fusion (Rackauckas, 2024) by utilizing reciprocal rank fusion of relevance and word frequency ranking to reach strong and robust retrieval performance. We introduce previous Agent-based works CoA (Zhang et al., 2024d) and LongAgent (Zhao et al., 2024a) to enhance the comprehensiveness of our baseline to highlight the improvement of our works.
Metrics. We evaluate the QA tasks and Misc by F1 Score (Bai et al., 2024) and Sequence Match Ratio [^5] https://docs.python.org/3/library/difflib.html for Retrieval, code tasks by exact match. Besides, we introduce fine-grained Progress Score $f$ from AgentBoard (Ma et al., 2024) for quantification of the multi-turn LLM agents’ performance which is denoted by:
4.2 Overall Results of XpandA
@@
Code Tasks, Retrieval & Misc. Table 2 displays the test results for these three classification tasks. Similarly, XpandA achieved optimal scores on the evaluation metrics for the majority of these tasks. However, it is noteworthy that RAG + Llama3.1-8B-Instruct delivered the best performance on the GraphWalks dataset, with XpandA ranking second. This outcome may be attributed to the fact that the text in GraphWalks is presented in a structured ”node-¿node” format (where each ’node’ is a hexadecimal hash string). Such a structure potentially allows RAG to retrieve the intended textual information with greater precision compared to attention-based language models.